# On The Topology of Polygonal Meshes

### v1

@@ 2.2 Intrinsic and Extrinsic Topology
The polygonal mesh representation encodes both the geometry and topology of shapes. While that is powerful, it can also be slightly confusing because we can consider the intrinsic topology of a polygonal mesh or its extrinsic topology of the mesh as a set of points in an embedding space. This leads to two different points of view, and it is necessary to choose a point of view before analysing the topology of a mesh.
The Extrinsic View
Extrinsically, a polygon is a set of points in space, $\bigcup_{f}P_{f}$, where $P_{f}$ is the set of points that belong to the polygonal face $f\in\mathcal{F}$. Thus, $P_{f}$ is only defined if we know the positions of the corners of $f$ and if we know how the geometry of the polygon is defined from those corners. In practice, polygons are often assumed to be almost flat and the question of how precisely the geometry is defined is ignored. As long as we stick to triangle meshes (which is often the case) this is unproblematic since a triangle is always flat. According to the extrinsic view, the polygonal mesh representation can be understood as a generalization of the (ancient) notion of a polyhedron. For instance, a cube is a polygonal mesh with eight vertices, twelve edges, and six faces. However, polyhedra usually have flat faces which is not required for polygonal meshes. We could twist the top face of the cube (relative to the bottom) and that would make the side faces non-planar. Most geometers would no longer consider it a polyhedron, but it is still a perfectly valid polygonal mesh. We conclude that a polyhedron is a polygonal mesh, but the converse is not necessarily true.
The Intrinsic View
From the intrinsic point of view, the positions of the vertices do not matter since the geometry is not important; a mesh is simply a schematic diagram of the structure of a surface. We can use this schematic to realise the mesh in 3D by assigning positions to the vertices. Edges are then typically realised as line segments connecting the vertices, and the faces become polygonal patches. However, the intrinsic topology is independent of the realisation. Expressed differently, we can see a polygonal mesh as a graph embedded in an abstract surface. Construed as a graph, some cycles of edges bound a patch on the manifold, and the faces of the polygonal mesh identify the cycles that bound patches.
In this document, we are most concerned with the intrinsic topology. Nevertheless, it is important to be aware of this distinction since the extrinsic and the intrinsic topology do not have to agree. Specifically, $\mathcal{M}$ may intersect itself in the embedding space which would often cause its extrinsic topology to differ from the intrinsic.
@@ 2.3 Qualitative Topology
The three topological invariants
  • manifoldness,
  • watertightness, and
  • orientability
are often important in the context of polygonal meshes. We can tell that they are topological properties since they are unaffected by a deformation of the mesh, but unlike the number of holes they are either true of false for a given mesh and thus qualitative rather than quantitative.
2.3.1 Manifoldness
  • Every edge of $\mathcal{M}$ is adjacent to either one or two faces, i.e. Eq. (3) is fulfilled.
  • The faces incident on a vertex form only a single connected component. We consider faces connected if they share an edge.
Another way of stating the vertex condition is that the union of the faces incident on the vertex should remain a single connected set if the vertex were removed. If $\mathcal{M}$ passes both of these tests, we know that it can everywhere be flattened to a disc and is thus intrinsically manifold.
Provided a mesh is manifold we can also prove that boundary edges always form cycles. As stated above, the faces incident on a vertex must form a single component. This means that a vertex is always incident on either two or zero boundary edges since the incident faces must form either a fan or a disc around the vertex. margin: [Uncaptioned image] All vertices which belong to boundaries are obviously incident on two boundary edges (see the figure on the right). This means that we can alway jump from a boundary edge to the next edge across one of its vertices, but, assuming the mesh is finite, we must eventually revisit the edge we started from.
A mesh which is produced by joining two pyramids (one upside down) such that the tips of the pyramids are a shared vertex is not manifold. We could try to flatten a region around the shared vertex, but it would always end up as two discs connected at a single point. Likewise, if two cubes are placed side by side such that two edges touch (forming a shared edge) that is also a non-manifold configuration since four faces then meet at the edge where the cubes are joined. Of course, we are here assuming that the shared vertex or edge is actually shared in the representation of the polygonal mesh. Otherwise, $\mathcal{M}$ remains manifold in the intrinsic sense though not in the embedding. In practice, manifoldness is a property which can easily be true intrinsically but not extrinsically due to self-intersections. The two non-manifold configurations are shown in Figure 3.
@@ 2.3.2 Watertightness
margin: [Uncaptioned image]
$\mathcal{M}$ is watertight if it divides space into an exterior region, an interior region, and $\mathcal{M}$ itself. If this is true, any path from a point in the interior to a point in the exterior contains at least one point that belongs to $\mathcal{M}$. A cube (and polyhedra in general) is watertight, but remove one of its faces (as shown in the figure on the right), and the property is lost. Again, this is the extrinsic definition of watertightness. Intrinsically, $\mathcal{M}$ is watertight if it is closed, i.e. it has no boundary. A mesh which is not watertight has at least one cycle of boundary edges.
@@ 2.3.3 Orientability
There are several famous examples of non-orientable surfaces such as the Möbius band and the Klein bottle, but in the following we will generally assume that the polygonal meshes we consider are orientable. Non-orientable surfaces are confusing in the context of the topology of polygonal meshes. It is relatively clear that a Möbius band is not homeomorphic to an annulus, so our strongest tool can distinguish this non-orientable surface from its closest orientable relative, but they have the same homotopy type, and the quantitative methods that we will discuss later cannot make this distinction: we get the same numbers for an annulus and a Möbius band. On the other hand, it is easy to detect that a surface is non-orientable. We simply have to assign consistent orientations to all its faces and realize that this is impossible. In fact, some polygonal mesh representations (in the computer science sense) cannot represent non-orientable meshes, and a seam will be introduced somewhere.
@@ 3 Euler’s Formula
which holds for any polygonal mesh that is a topological sphere. A polygonal mesh is considered to be a topological sphere if it is homeomorphic to a sphere. Some of the most famous examples of polygonal meshes which are also topological spheres are the Platonic solids which are shown in Figure 4. By consulting Table 1 we can ascertain that they all fulfil Eq. 4.
@@
margin: [Uncaptioned image] This mesh is a topological sphere.
@@
Figure 5: A simplified version of the Stanford Armadillo is shown left. On the right is a mesh consisting of the exact same vertices and faces but transformed such that all vertices lie on the sphere. The edges remain straight lines, but it is east to imagine that they could be replaced with great arcs: if we imagine a light placed in the centre of the sphere, the shadows that the edges would cast on the (inside of) the sphere would be precisely the needed circular arcs as explained by Eppstein in his Geometry Junkyard.
@@
where $k_{f}$ is the number of sides of $f$ and $\alpha_{i}$ is the angle of corner $i$ of the polygon. See Tristan Needham [needham2021visual] for a proof of this formula for spherical triangles. The generalization to spherical polygons is straight forward since polygons can always be triangulated. We can plug (6) into (5), obtaining
$\sum_{f=1}^{\mathcal{F}}\left(\sum_{i=1}^{k_{f}}\alpha_{i}-(k_{f}-2)\pi\right)=4\pi$
@@ 4 Quantitative Topology
None of these three properties can be obtained directly from Euler’s formula. Fortunately, there is a generalization known as the Euler-Poincaré formula:
@@
$s$ is easy to interpret. It is simply the number of connected components of the mesh. As, we recall there is no notion of solids when we deal with polygonal meshes. Hence, we do not distinguish between mesh components that represent interior voids and separate shapes even if it may be easy to tell these cases apart based on the geometry. If the mesh does have sphere topology, $s=1$, $g=0$, and $b=0$, and we immediately see that the formula reduces to the familiar one in this case. The generalization to $s>1$ is trivial. If we have two objects of sphere topology and add up their vertices, edges, and faces, we get
VE+F=2s.V-E+F=2s\,.
@@
$V-E+F^{\prime}=2-1$
where $F^{\prime}=F-1$ since we removed one face. In other words, removing a single face just amounts to subtracting one from the right-hand side. What if the hole is bigger than a single face? We soon realise that removing more than one face does not change the left-hand side: when one more faces is removed to enlarge a hole, we also remove the edge that separates this faces from the hole. In some cases a face shares more than one edge with a hole, but then a vertex is also removed, and the end result would always be that the left-hand side is not changed further. We arrive at
VE+F=2sbV-E+F=2s-b
@@
Arguably, $g$, the genus of the shape, is the “poster entity” of topological invariants, and it is equal to the number of handles of the shape. A handle is the same as a hole, since we can always deform (and homeomorphically map) a shape with $N$ holes into a sphere with $N$ handles. Thus, $g=1$ for a torus which has a single hole/handle and $g=0$ for a sphere that has none.
@@
$V^{\prime}-E^{\prime}+F^{\prime}=(V+2N)-(E+3N)+(F+N)=V-E+F\,.$
@@
VE+F=2s2gb.V-E+F=2s-2g-b\,.
Thus, the general Euler-Poincaré formula has been proven. To be fair, our proof does not meet the standards of rigor you would expect from a text aimed at mathematicians, but leaving that aside we can be certain that Eq. (7) holds.
@@ 5 The Betti Numbers
Having proven that the Euler-Poincaré formula is true does not seem to have been very helpful for understanding it. However, there is a clue in its structure: it looks like the beginning of an alternating sum since the first term, $V$, is added, the next term, $E$, is subtracted, and the final term, $F$, is again added. We will now show that VE+FV-E+F is equal to another alternating sum which has a much more clear interpretation. To do so, we need to make the notion of cycles precise.
5.1 Cycles
Cycles are sets of mesh entities which have no boundary. While the name evokes “cycles of edges”, we will also define vertex cycles and face cycles.
Vertex Cycles
Any single vertex is a cycle. This is simply a matter of definition: we define a vertex to have no boundary. Consequently, a vertex is a cycle. Any collection of vertices is also a cycle.
Edge Cycles
Loosely speaking, an edge cycle is a set of edges that “bites its own tail”. Eq. (2) lays down the more precise condition that the edges of a face form a cycle if any vertex of the mesh belongs to either zero or two edges of the face. Note that we can make more complex cycles by combining such simple cycles and in this case a vertex can be included more than two times, but always an even number of times.
Face Cycles
A set of faces is a cycle if every edge of every face is adjacent to precisely two faces. In other words, a set of faces is a cycle if it is closed (or watertight). Again, a collection of face cycles is still a cycle.
Please refer to Figure 7 for an illustration of edge cycles and face cycles.
Figure 7: Left: two sets of edges where the left one is not a cycle, but the right one is a cycle. The cycle instigator is shown as a red line segment. Right: a cube which is not a cycle since it is missing a face is shown on the left. On the right a closed cube which constitutes a cycle.
@@ 5.2 Cycle Instigators
The next task is to count how many vertices, edges, and faces, we can choose without creating any cycles. To this end, we will introduce the notion of a cycle instigator. Given a set of mesh edges, a cycle instigator is an edge, $e^{i}$, that becomes part of a cycle when added to the set. To be precise, $e^{i}$ is a cycle instigator with respect to a set of edges $\mathcal{T}$ iff
$\exists\mathcal{C}:\mathcal{C}\subseteq\mathcal{T}\cup e^{i}\wedge e^{i}\in\mathcal{C}$
@@
In fact, this question is easy to answer. Assuming $\mathcal{M}$ is a single component, a spanning tree of the vertices of $\mathcal{M}$ is a collection of edges which is incident on every vertex of $\mathcal{V}$. Now, a tree contains no cycles by definition, and we can create a tree by first adding any edge of the mesh to $\mathcal{T}$ and subsequently adding only edges that connect a vertex already incident on $\mathcal{T}$ to a vertex on which $\mathcal{T}$ is not incident. If we count the edges while adding them, it is clear that our count reaches V1V-1 since we add one edge for every vertex except the first. To answer the question, we now note that any edge that does not belong to the tree defines a unique cycle and is thus an instigator. We can obtain this cycle by adjoining the instigating edge to the path we obtain by following the tree edges from its end points back towards the root until they meet at a vertex (which may be the root vertex). This procedure is shown in Figure 8. We conclude that the maximal number of non-instigating edges is the number of edges in a spanning tree, i.e. V1V-1, and we can now define the number of non-instigating edges, $E_{N}$, and the number of cycle-instigating edges, $E_{C}$, as follows:
@@ 5.3 Reformulating Euler’s Formula
Likewise, when we add an instigating edge, it is not a single cycle but a subspace in a space of edge cycles that is created. This is because the instigator defines a unique cycle that can be construed as a basis vector. If we later add a non-instigating face that fills in the cycle, the boundary of that face also represents a basis vector which we use to project out a dimension of the space, decreasing $\beta_{1}$ by 1. A geometric way to think about it is that adding the face allows the curve to “slip across it” making edge cycles indistinguishable regardless of which way they go around the face. The upshot is that $\beta_{1}$ represents the dimension of the space of edge cycles that is not matched by faces.
@@ 6 Examples
To find $s$, $g$, and $b$ we follow the procedure outlined in Section 4.0.1 and the Betti numbers are simply computed using Equations (18), (19), and (20). The quantitative properties of the four example shapes are summarized in Table 2.
@@
Of course, we must be a little careful. Imagine an object which is a single edge shared by three trianglesmargin: [Uncaptioned image] . This object consists of 5 vertices, 7 edges, and 3 faces. We cannot compute the number of boundary curves, since they do not form closed cycles. We can compute the Betti numbers, though, and get $\beta_{0}=1$, $\beta_{1}=0$, and $\beta_{2}=0$. In other words, while this object is non-manifold it has the Betti numbers of a topological disc. Likewise, we could turn the annulus into a Möbius band and the torus into a Klein bottle – all without changing their respective Betti numbers since these objects have the same numbers of vertices, edges, and faces.
@@
7 Turning Polygonal Meshes into Discs
@@
For shapes of higher genus, i.e. $g>0$, it turns out that we need $2g$ closed cuts. Recall that for an object that consists of a single component and which has no boundary curves $\beta_{1}=2g$ and, for simplicity, we will assume that the object has no boundary curves. We will also assume there is just a single connected component. This means that the number of cuts we need to make is exactly the number of unique edge cycles which do not bound sets of faces. It turns out that we also need to make the required cuts exactly along these edge cycles.
In the following, we will present an algorithm originally due to Eppstein [eppstein2003dynamic], improved by Erickson and Whittlesey [erickson2005greedy] and discussed by Keenan Crane [crane_loops].
Figure 10: Left: A torus which has been cut and flattened to a disc. Note that the edges on the top are identical to those on the bottom and likewise, the left and right edges are identical. The edges of the shortest-path tree are shown pale blue, and the co-tree edges are indicated by bold arrows. In both cases, edges point from parent to child. The root of the shortest-path tree is the top left corner, and the root of the co-tree is the face indicated by a yellow star. The edges highlighted by red and purple are the cycle-instigating edges. Right: a rendering of the torus as a 3D object with the image on the left shown as a texture.
@@
The first step is to create a spanning tree of the vertices. We will do this using Dijkstra’s algorithm [cormen2022introduction] which produces a tree of shortest paths from an initial vertex, $v_{0}$: the path from any vertex $v$ to $v_{0}$ along the tree is a shortest path between $v$ and $v_{0}$. Just like any other spanning tree, the shortest-path tree contains $E_{N}=V-1$ edges. A shortest-path tree computed for a torus is shown in Figure 10.
@@
The next step is to construct a so-called co-tree. The co-tree is built from the dual of the mesh: a graph whose vertices are the faces of the mesh and two dual vertices are connected if the faces share an edge. We can build a co-tree using an algorithm that is similar to Dijkstra’s. The key component is a queue that keeps track of faces which have not yet been visited but are neighbours of visited faces.
An arbitrary starting face, $f_{0}$, is added to a queue. In each iteration, we pick a face $f$ from the queue. If $f$ is not in the set of visited faces, we add it to this set and then we consider each of its neighbors, i.e. each face, $f_{n}$, which shares an edge with $f$. If the edge shared by $f$ and $f_{n}$ does not belong to the shortest-path tree, we add $f_{n}$ to the queue and mark the shared edge as visited. This process is iterated until the queue is empty. A co-tree for a torus is shown in Figure 10.
@@
Simplistic Cutting Procedure
Assume now that we were to cut every edge in the shortest-path tree. The result would not have disc topology yet, but if we were to also cut the remaining $\beta_{1}$ edges, we would be left with a mesh where the only uncut edges were those that belong to the co-tree. Since a tree is connected and has no loops, this mesh would have disc topology.
Actual Cutting Procedure
However, we do not have to cut every edge of the shortest-path tree. What we do instead is to trace back from the end points of each of the $\beta_{1}$ cycle-instigating edges along the edges of the shortest-path tree. That produces $\beta_{1}$ cycles. Cutting along these edge cycles cannot disconnect the faces of the mesh since we are cutting along a subset of the edges of the simplistic procedure discussed above. However, there are no more cycles left which do not bound regions of the mesh. Consequently, these cuts must produce a mesh of disc topology. For a torus, $\beta_{1}=2$ and the two instigator edges are shown in purple and red in Figure 10. Note that in case of the red edge, the right end point is already connected to the root of the shortest path tree; in order to close the loop, we just have to trace back along the horizontal edges to the root vertex (in the top left corner). Likewise, to close the loop defined by the purple instigator, we simply trace vertically from the top vertex of the instigator to the root. It is clear that cutting along these two curves will turn the torus into a rectangle. Note that the top and bottom edges are really the same edges, and the same is true of the right side and left side edges.
Borrowing a term from homology, we will refer to the edge-cycles as generator curves. This is mostly because we need an ambiguous term to describe the curves that must be cut in order to turn the mesh into one that has disc topology. They are generators, though, in the sense that variations of these curves can be used to generate the entire group $H_{1}$.
7.1 Example
With the algorithm just described, we can compute generator curves and cut the mesh along these curves, turning it into a topological disc.
@@ 7.1 Example
Figure 11 shows the Fertility model together with a starting vertex (orange) and eight pictures that show the individual generator curves. In Figure 12 we see the Fertility model being flattened into a disc after it has been cut along a set of generator curves

### v2

@@ 2.5 Intrinsic and Extrinsic Topology
We have to face the fact that the topology of a polygonal mesh can mean two different things.
  • If we are interested in the intrinsic topology, the geometry is not important; a mesh is simply a schematic diagram of the structure of a surface. We can use this schematic to realise the mesh in 3D, but the intrinsic topology is independent of the realisation.
  • If we study the extrinsic topology, the realisation is the polygonal mesh. According to this view, a polygonal mesh is simply a set of points in 3D, namely the union of the points in contained in the faces. The topology of the mesh is that of the point set.
It is important to be aware of this distinction since the extrinsic and the intrinsic topology do not have to agree. Specifically, $\mathcal{M}$ may intersect itself in the embedding space which would often cause its extrinsic topology to differ from the intrinsic. For instance, a mesh may be intrinsically manifold but two vertices or two edges may touch as shown in Figure 5. In this case, the mesh is extrinsically non-manifold despite being manifold from the intrinsic point of view.
@@ 2.6 Qualitative Topology
The topological invariants watertightness and orientality are often important in the context of polygonal meshes. We can tell that they are topological properties since they are unaffected by a deformation of the mesh, but unlike the number of holes they are either true of false for a given mesh and thus qualitative rather than quantitative.
@@ 2.6.1 Watertightness
$\mathcal{M}$ is watertight if it divides space into an exterior region, an interior region, and $\mathcal{M}$ itself. If this is true, any path from a point in the interior to a point in the exterior contains at least one point that belongs to $\mathcal{M}$. A cube (and polyhedra in general) is watertight, but remove one of its faces (as shown in the figure on the right), and the property is lost. margin: [Uncaptioned image]
Not watertight.
Again, this is the extrinsic definition of watertightness. Intrinsically, $\mathcal{M}$ is watertight if it is closed, i.e. it has no boundary. A mesh which is not watertight has at least one cycle (cf. Section 2.2) of boundary edges.
@@ 2.6.2 Orientability
There are several famous examples of non-orientable surfaces such as the Möbius band and the Klein bottle, but in the following we will generally assume that the polygonal meshes we consider are orientable. Non-orientable surfaces are confusing in the context of the topology of polygonal meshes. It is relatively clear that a Möbius band is not homeomorphic to an annulus, so our strongest tool can distinguish this non-orientable surface from its closest orientable relative, but they have the same homotopy type, and the quantitative methods that we will discuss later cannot make this distinction: we get the same numbers for an annulus and a Möbius band. On the other hand, it is easy to detect that a surface is non-orientable. We simply have to assign consistent orientations to all its faces and realise that this is impossible. In fact, some polygonal mesh representations (in the computer science sense) cannot represent non-orientable meshes, and a seam will be introduced somewhere.
@@ 3 Euler’s Formula
which holds for any polygonal mesh that is a topological sphere. margin: [Uncaptioned image] This mesh is a topological sphere. A polygonal mesh is considered to be a topological sphere if it is homeomorphic to a sphere. Some of the most famous examples of polygonal meshes which are also topological spheres are the Platonic solids which are shown in Figure 6. By consulting Table 1 we can ascertain that they all fulfil Eq. 14.
@@
Figure 7: A simplified version of the Stanford Armadillo is shown left. On the right is a mesh consisting of the exact same vertices and faces but transformed such that all vertices lie on the sphere. The edges remain straight lines, but it is easy to imagine that they could be replaced with great arcs: if we imagine a light placed in the centre of the sphere, the shadows that the edges would cast on the (inside of) the sphere would be precisely the needed circular arcs as explained by Eppstein in his Geometry Junkyard.
@@
where $k_{f}$ is the number of sides of $f$ and $\alpha_{i}$ is the angle of corner $i$ of the polygon. See Tristan Needham [needham2021visual] for a proof of this formula for spherical triangles. The generalisation to spherical polygons is straightforward since polygons can always be triangulated. We can plug (16) into (15), obtaining
$\sum_{f=1}^{\mathcal{F}}\left(\sum_{i=1}^{k_{f}}\alpha_{i}-(k_{f}-2)\pi\right)=4\pi$ (17)
@@ 4 Quantitative Topology
None of these three properties can be obtained directly from Euler’s formula. Fortunately, there is a generalisation known as the Euler-Poincaré formula:
@@
$s$ is easy to interpret. It is simply the number of connected components of the mesh. As, we recall there is no notion of solids when we deal with polygonal meshes. Hence, we do not distinguish between mesh components that represent interior voids and separate shapes even if it may be easy to tell these cases apart based on the geometry. If the mesh does have sphere topology, $s=1$, $g=0$, and $b=0$, and we immediately see that the formula reduces to the familiar one in this case. The generalisation to $s>1$ is trivial. If we have two objects of sphere topology and add up their vertices, edges, and faces, we get
VE+F=2s.V-E+F=2s\,. (21)
@@
$V-E+F^{\prime}=2-1$ (22)
where $F^{\prime}=F-1$ since we removed one face. In other words, removing a single face just amounts to subtracting one from the right-hand side. What if the hole is bigger than a single face? We soon realise that removing more than one face does not change the left-hand side: when a face is removed to enlarge a hole, we also remove the edge that separates this faces from the hole. In some cases a face shares more than one edge with a hole, but then a vertex is also removed, and the end result would always be that the left-hand side is not changed further. We arrive at
VE+F=2sb.V-E+F=2s-b\,. (23)
@@
Arguably, $g$, the genus of the shape, is the “poster entity” of topological invariants, and it is equal to the number of handles of the shape. A handle is the same as a hole (in the donut sense), since we can always deform (and homeomorphically map) a shape with $N$ holes into a sphere with $N$ handles. Thus, $g=1$ for a torus which has a single handle and $g=0$ for a sphere that has none.
@@
$V^{\prime}-E^{\prime}+F^{\prime}=(V+2N)-(E+3N)+(F+N)=V-E+F\,.$ (24)
@@
VE+F=2s2gb.V-E+F=2s-2g-b\,. (25)
Thus, the general Euler-Poincaré formula has been proven – or at least the proof has been sketched since we glossed over some details.
@@ 5 The Betti Numbers
Having proven that the Euler-Poincaré formula is true does not seem to have been very helpful for understanding it. However, there is a clue in its structure: it looks like the beginning of an alternating sum since the first term, $V$, is added, the next term, $E$, is subtracted, and the final term, $F$, is again added. We will now show that VE+FV-E+F is equal to another alternating sum which has a much more clear interpretation.
@@ 5.1 Cycle Instigators
The first task is to count how many vertices, edges, and faces, we can choose without creating any cycles. To this end, we will introduce the notion of a cycle instigator. Given a set of mesh edges, a cycle instigator is an edge, $e^{i}$, that becomes part of a cycle when added to the set. To be precise, $e^{i}$ is a cycle instigator with respect to a set of edges $\mathcal{T}$ iff
$\exists\mathcal{C}:\mathcal{C}\subseteq\mathcal{T}\cup e^{i}\wedge e^{i}\in\mathcal{C}$ (26)
@@
In fact, this question is easy to answer. Assuming $\mathcal{M}$ is a single component, a spanning tree of the vertices of $\mathcal{M}$ is a collection of edges which is incident on every vertex of $\mathcal{V}$. Now, a tree contains no cycles by definition, and we can create a tree by first adding any edge of the mesh to $\mathcal{T}$ and subsequently adding only edges that connect a vertex already incident on $\mathcal{T}$ to a vertex on which $\mathcal{T}$ is not incident. If we count the edges while adding them, it is clear that our count reaches V1V-1 since we add one edge for every vertex except the first. To answer the question, we now note that any edge that does not belong to the tree defines a unique cycle and is thus an instigator. We can obtain the instigated cycle by adjoining the instigating edge to the path we obtain by following the tree edges from its end points back towards the root until they meet at a vertex (which may be the root vertex). This procedure is shown in Figure 9. We conclude that the maximal number of non-instigating edges is the number of edges in a spanning tree, i.e. V1V-1, and we can now define the number of non-instigating edges, $E_{N}$, and the number of cycle-instigating edges, $E_{C}$, as follows:
@@ 5.2 Reformulating Euler’s Formula
Likewise, when we add an instigating edge, it is not a single cycle but a subspace in a space of edge cycles that is created. This is because the instigator defines a unique cycle that can be construed as a basis vector. If we later add a non-instigating face that fills in the cycle, the boundary of that face also represents a basis vector which we use to project out a dimension of the space, decreasing $\beta_{1}$ by 1. A geometric way to think about it is that adding the face allows the curve to “slip across it” making edge cycles indistinguishable regardless of which way they go around the face. The upshot is that $\beta_{1}$ represents the dimension of the space of edge cycles that do not bound faces.
@@ 6 Examples
To find $s$, $g$, and $b$ we follow the procedure outlined in Section 4.1 and the Betti numbers are simply computed using Equations (39), (40), and (41). The quantitative properties of the four example shapes are summarised in Table 2.
margin: [Uncaptioned image]
@@
Of course, we must be a little careful. Imagine an object which is a single edge shared by three triangles (shown right). This object consists of 5 vertices, 7 edges, and 3 faces. We cannot compute the number of boundary curves, since they do not form closed cycles. We can compute the Betti numbers, though, and get $\beta_{0}=1$, $\beta_{1}=0$, and $\beta_{2}=0$. In other words, while this object is non-manifold it has the Betti numbers of a topological disc. Likewise, we could turn the annulus into a Möbius band and the torus into a Klein bottle – all without changing their respective Betti numbers since these objects have the same numbers of vertices, edges, and faces.
@@
This informs us that the quantitative topological properties should be understood in light of the qualitative topological properties. It is not enough to know $V$, $E$, and $F$. We also need to know if the polygonal mesh is manifold, whether it has a boundary, and if it is orientable – in order to meaningfully compute the quantitative topological properties.
@@
7 Computing a Cut Graph
@@
For shapes of higher genus, i.e. $g>0$, we need $2g$ closed cuts. Recall that for an object which consists of a single component and which has no boundary curves $\beta_{1}=2g$ and, for simplicity, we will assume that the object has no boundary curves. We will also assume there is just a single connected component. This means that the number of cuts we need to make is exactly the number of unique edge cycles which do not bound sets of faces.
In the following, we will present an algorithm originally due to Eppstein [eppstein2003dynamic] and improved by Erickson and Whittlesey [erickson2005greedy]. Later Keenan Crane presented the algorithm with some modifications [crane_loops].
Figure 11: Left: A torus which has been cut and flattened to a disc. Note that we should consider the bottom to be attached to the top and likewise the right side is attached to the left. The edges of the shortest-path tree ($\mathcal{T}$) are shown pale blue, and the co-tree edges ($\mathcal{T}^{c}$) are those which are crossed by the bold green arrows. Edges point from parent to child in both trees, and the root of the shortest-path tree is the orange disc, and the root of the co-tree is the yellow star. The two edges highlighted by red and purple, respectively, are the cycle-instigating edges. The $\Gamma$ shaped frame highlights the edges in $\mathcal{B}$. Right: a rendering of the torus as a 3D object with the image on the left shown as a texture.
@@
Given a mesh, $\mathcal{M}$, the first step is to create a spanning tree, $\mathcal{T}\subset\mathcal{E}$, of the vertices. We can do this using Dijkstra’s algorithm [cormen2022introduction] which produces a tree of shortest paths from the root vertex to any vertex of the mesh. Just like any other spanning tree, the shortest-path tree contains $E_{N}=V-1$ edges. A shortest-path tree computed for a torus is shown in Figure 11.
@@
The next step is to construct a so-called co-tree, $\mathcal{T}^{c}\subset\mathcal{E}$, which is built from the dual of the mesh: a graph whose vertices are the faces of the mesh and two dual vertices are connected if the faces share an edge. We can build a co-tree using an algorithm that is similar to Dijkstra’s. The key component is a queue that keeps track of faces which have not yet been visited but are neighbours of visited faces.
[…]
@@
Constructing the Cut Graph
The cut graph is used to turn the mesh into one that has disc topology. However, there is actually a simple procedure that produces a mesh of disc topology, $\mathcal{M}^{\prime}$, without any need for the cut graph. This procedure is to cut along every edge except those that belong to $\mathcal{T}^{c}$. Since the co-tree is connected and has no loops, $\mathcal{M}^{\prime}$ has disc topology.
@@
We now claim that cutting $\mathcal{M}$ along the cut graph, $\mathcal{B}$, turns it into a topological disc which we will denote $\mathcal{M}^{*}$.
To prove this, observe that if we were to cut $\mathcal{M}^{*}$ along all not-already-cut edges that belong to $\mathcal{T}$, i.e. edges in $\mathcal{T}\backslash\mathcal{B}$, we arrive at $\mathcal{M}^{\prime}$ which is known to have disc topology. Since $\mathcal{T}\backslash\mathcal{B}\subset\mathcal{T}$, it does not contain any cycles. In other words, we can transform $\mathcal{M}^{*}$ into $\mathcal{M}^{\prime}$ without additional cyclic cuts. Hence $\mathcal{M}^{*}$ must be a topological disc – possibly with additional boundary curves.
We state without proof that if the cut graph consists of a single connected component, a disc topology mesh is produced; otherwise a disc with n1n-1 holes where $n$ is the number of connected components of the cut graph.
7.1 Connections to Algebraic Topology
The edge-cycles that belong to the cut graph are known as generators of the classes in the first homology group. By adding boundaries of collections of faces to generators and combinations of generators, it is possible to obtain any edge cycle on a mesh. The generators produced by the described algorithm are also sometimes called a homotopy basis, e.g. by Keenan Crane [crane_loops]. Grossly simplifying, the difference between a homotopy basis and a homology basis is that the generators of a homotopy basis have a common base point (the root of our shortest-path tree). For a homology basis we could choose different generator curves which do not have the same starting point. This can be advantageous if we care about the geometry of the generators. For instance, the HanTun Algorithm by Dey et al. [dey2008computing] finds handle and tunnel loops which together form a homology basis. Note, though, that in this case the generators of the homology basis may not form a connected cut graph.
7.2 Examples
With the algorithm just described, we can compute a cut graph that turns a closed mesh of any genus into a topological disc.
For a torus, $\beta_{1}=2$ and the two instigator edges are shown in purple and red in Figure 11. Note that in case of the red edge, the right end point is already connected to the root of the shortest path tree; in order to close the loop, we just have to trace back along the horizontal edges to the root vertex (in the top left corner). Likewise, to close the loop defined by the purple instigator, we simply trace vertically from the top vertex of the instigator to the root. It is clear that cutting along these two curves will turn the torus into a rectangle. Note that the top and bottom edges are really the same edges, and the same is true of the right side and left side edges.
[…]
Figure 12 shows the Fertility model together with a starting vertex (orange) and eight pictures that show the individual generator curves from the cut graph. In Figure 13 we see the Fertility model being flattened into a disc after it has been cut along the cut graph.